Solve the equation is 2A/3=8+4A
Answer by stanbon(75887)

2A/3=8+4A
---
Multiply both sides by 3 to get:
2A = 24+12A
Subtract 12A from both sides to get:
-10A = 24
A = -12/5
-----------
Cheers,
Stan H.